A mason can lay the bricks in a sidewalk in 12 hours. The mason's apprentice requires 16 hours to do the same job. After working together for 4 hours, the mason leaves for another job and the apprentice continues working. How long will it take the apprentice to complete the job?
Answer by stanbon(75887)

A mason can lay the bricks in a sidewalk in 12 hours. The mason's apprentice requires 16 hours to do the same job. After working together for 4 hours, the mason leaves for another job and the apprentice continues working. How long will it take the apprentice to complete the job?
-----------
Mason DATA:
Time= 12 hr/job ; Rate = 1/12 job/hr
---------------
Aprentice DATA:
Time = 16 hr/job; Rate = 1/16 job/hr
-----------------
Togeether DATA:
Rate = (1/12+1/16) = 28/(12*16)= 7/48 job/hr
If they work together for 4 hrs they with do 4(7/48)=7/12 of the job.
---------------
EQUATION:
Let the amount of time the apprentice takes to complete the job be x hrs
xhr(1/16)job/hr=(1-7/12)job
x/16 = 5/12
x=20/3 hrs
x=6.33333... hrs
The apprentice will take 6 1/3 hrs to complete the job
=======================
Cheers,
Stan H.
--------------